electrons are emitted with zero velocity from a certain metal surface when it is exposed to radiations of wavelength 7000 A0​ . The work function of the metal is ?

Grade:12th pass

2 Answers

Arun
4 years ago
c = νλ∴ νo = c/λo = 3.0×108 ms-1/7000×10-10 m = 4.29×1014 s-1Work function (wo) = hνo = 6.626×10-34 Js×4.29×1014 s-1 = 2.842×10-19 J 
Vikas TU
4 years ago
E=hc/lambda. 
also E=hvo
C/lambda= vo
3 x 10^8 / 7000 x 10^-10 .       ( 1 A = 10-10)
vo = 4.281 x 10^14  s^-1
w0 = hv0 = 6.63* 10^-34 * 4.28*10^14
=> 28.41 *10^-20 J